### 17 Mar 2026 v2

# Investigating $\Omega_{c}$ spectroscopy in two-body $\Omega_{b}$ decays

Abstract
We investigate the $1S$-, $1P$-, $2S$-, and $1D$-wave $\Omega_{c}(css)$ spectroscopy through the non-leptonic decays of $\Omega_{b}^{-}$ baryon within the constituent quark model. For the lowest-lying $1S$-wave $\Omega_{c}$ state, we obtain the branching fractions ${\cal B}(\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-},\Omega_{c}^{0}\rho^{-})=(1.2,6.3)\times 10^{-3}$, which are consistent with existing model predictions. For $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$, and $\Omega_{c}(3090)$, observed in the proton-proton and $e^{+}e^{-}$ collisions and interpreted as members of the $1P$-wave multiplet (collectively denoted as $\Omega_{c}^{**}$), we predict the branching fractions of $\Omega_{b}^{-}\to\Omega_{c}^{**}\pi^{-},\Omega_{c}^{**}\rho^{-}$ at the level of $10^{-3}$. Assigning the newly observed $\Omega_{c}(3327)$ baryon to a $1D$-wave excitation with $J^{P}=5/2^{+}$ or $7/2^{+}$, we obtain ${\cal B}[\Omega_{b}^{-}\to\Omega_{c}(3327)^{0}\pi^{-},\Omega_{c}(3327)^{0}\rho^{-}]=(2.0,5.7)\times 10^{-3}$ or $(4.6,0.8)\times 10^{-3}$, respectively. The pronounced differences between these two scenarios provide a clear discriminant that can be tested in future measurements at LHCb.
I Introduction
A singly charmed $\Omega_{c}$ baryon is composed of three quarks, csscss, in which the two strange quarks form a correlated diquark subsystem. The orbital motion between the charm quark and the light diquark plays a central role in shaping the $\Omega_{c}$ spectrum, while possible internal orbital excitations of the diquark may also contribute. In addition, the spin of the ssss diquark and the spin of the charm quark provide further degrees of freedom, whose couplings with the orbital angular momentum give rise to the observed pattern of spin-parity states. Within the framework of heavy-quark effective theory, a systematic classification of the $\Omega_{c}$ spectrum has been established  Chen:2022asf ; Roberts:2007ni ; Yoshida:2015tia ; Luo:2025sns . Experimental measurements therefore play a key role in validating and refining the established $\Omega_{c}$ spectroscopy.
Several $\Omega_{c}$ baryons have been observed in the proton-proton collision by LHCb LHCb:2017uwr ; LHCb:2023sxp and the $e^{+}e^{-}$ collision by Belle Belle:2017ext . However, not all of the observed states have been unambiguously assigned definite quantum numbers LHCb:2021ptx . For instance, the $\Omega_{c}(3327)$ baryon, newly observed in Ref. LHCb:2023sxp , has been interpreted as a $1D$ state in several theoretical studies Zhong:2025oti ; Wang:2023wii ; Luo:2023sra ; Pan:2023hwt ; Yu:2023bxn ; Garcia-Tecocoatzi:2024aqz , while its spin-parity quantum numbers $J^{P}$ remain undetermined. The excited $\Omega_{c}$ states observed in Refs. LHCb:2017uwr ; Belle:2017ext comprise five states in total and collectively denoted as $\Omega_{c}^{**}$ by LHCb LHCb:2017uwr . These states have been extensively investigated within a wide variety of theoretical frameworks  Kucukyilmaz:2025rsd ; Li:2025frt ; Weng:2024roa ; Li:2024zze ; Karliner:2017kfm ; Garcia-Tecocoatzi:2022zrf ; Karliner:2023okv ; Ortiz-Pacheco:2023kjn ; Bahtiyar:2020uuj ; Luo:2023sra ; Garcia-Tecocoatzi:2024aqz ; Pan:2023hwt ; Yu:2023bxn ; Yu:2022ymb ; Chen:2017sci ; Agaev:2017lip ; Oudichhya:2023awb ; Jakhad:2023mni ; Padmanath:2017lng ; Agaev:2017jyt ; Wang:2017zjw ; Aliev:2017led ; Agaev:2017lip ; Wang:2017xam ; Chen:2017sci ; Aliev:2018uby ; Cheng:2017ove ; Oudichhya:2021kop ; Oudichhya:2023awb ; Wang:2017hej ; Wang:2017kfr ; Yao:2018jmc ; Wang:2017vnc ; Ali:2017wsf ; Chen:2017gnu ; Zhao:2017fov ; Garcia-Tecocoatzi:2022zrf ; Kim:2017khv ; Santopinto:2018ljf ; Jia:2020vek ; Luo:2023sra . Among the $\Omega_{c}^{**}$ states, $\Omega_{c}(3000)$, $\Omega_{c}(3050,3065)$, $\Omega_{c}(3090)$ are commonly interpreted as members of the $1P$ multiplet with quantum numbers $J^{P}=1/2^{-}$, $3/2^{-}$, and $5/2^{-}$, respectively, while the quantum numbers of $\Omega_{c}(3120)$ remain uncertain. Alternatively, the $\Omega_{c}^{**}$ states have also been proposed as more exotic configurations, such as hadronic molecules or pentaquark states Ozdem:2024ydl ; Wang:2018alb ; Wang:2021cku ; Yang:2017rpg ; Huang:2017dwn ; An:2017lwg ; Anisovich:2017aqa ; Nieves:2017jjx ; Liu:2018bkx ; Chen:2017xat ; Montana:2018teh ; Montana:2017kjw ; Ramos:2020bgs ; Praszalowicz:2022hcp ; Kim:2017jpx ; Debastiani:2017ewu ; Debastiani:2018adr ; Wang:2017smo ; Huang:2018wgr ; Zhu:2022fyb ; Xin:2023gkf ; Yan:2023ttx ; Yan:2023tvl ; Ozdem:2023okg ; Feng:2023ixl . Clearly, further theoretical and experimental investigations are required to clarify the nature and quantum-number assignments of these excited $\Omega_{c}$ baryons.
The LHCb Collaboration has recently observed the resonant decays $\Omega_{b}^{-}\to\Omega_{c}^{**}\pi^{-},\Omega_{c}^{**}\to\Xi_{c}^{+}K^{-}$ LHCb:2021ptx , in which the states $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$, and $\Omega_{c}(3090)$ reappear as resonance structures in the $\Xi_{c}^{+}K^{-}$ invariant mass spectrum. As a result, the existence of these states, established in the pppp and $e^{+}e^{-}$ collisions, is independently confirmed in weak decays. This observation opens up a new avenue for investigating the $\Omega_{c}$ spectroscopy, while simultaneously raising several important questions. First, the $\Omega_{c}(3120)$ baryon remains the only $\Omega_{c}^{**}$ state that has not been observed in this decay mode. Second, the $\Omega_{c}(3327)$ state, as well as other higher excited $\Omega_{c}$ baryons, can in principle also be produced in $\Omega_{b}^{-}$ decays; however, no corresponding signals were reported until very recently. Third, although four $\Omega_{c}^{**}$ states have been re-established in weak decays, their branching fractions have not yet been measured. In view of these issues, we consider the decays $\Omega_{b}^{-}\to\Omega_{c}^{(*)0}M$ as a suitable test ground. These channels involve the lowest-lying $1S$ baryon $\Omega_{c}^{0}$, as well as the excited $1P$- and $1D$-wave states. The final-state meson $M$, where $M$ can be $\pi$ or $\rho$, provides scalar or vector quantum numbers that couple differently to $\Omega_{c}^{(*)0}$ states.
Theoretical studies of semileptonic and charmful $\Omega_{b}$ decays Ortiz-Pacheco:2020hmj ; Hsiao:2021mlp ; Wang:2024mjw ; Duan:2025aur , as well as the two-body processes $\Omega_{b}\to\Omega_{c}M$ Cheng:1996cs ; Ivanov:1997hi ; Ivanov:1997ra ; Gutsche:2018utw ; Zhao:2018zcb ; Neishabouri:2026mjn ; Patel:2025gbw , remain rather limited in the literature. In particular, the decays $\Omega_{b}\to\Omega_{c}^{*}M$ have been much less extensively explored Li:2021kfb ; Chua:2018lfa ; Chua:2019yqh . Motivated by this situation, we investigate the decays $\Omega_{b}^{-}\to\Omega_{c}^{(*)0}M$ within the non-relativistic constituent quark model (CQM). The CQM has proven to be a reliable framework in describing hadronic decays of hyperons Niu:2020aoz , $D$ mesons Cao:2023csx ; Cao:2023gfv , and heavy baryons containing $b$ or $c$ quarks Niu:2020gjw ; Wang:2022zja ; Niu:2021qcc ; Pervin:2006ie ; Pervin:2005ve . These processes therefore provide useful probes for exploring the $1S$-, $1P$-, $2S$-, and $1D$-wave structures in the $\Omega_{c}$ spectrum.
II Formalism
The decays $\Omega_{b}^{-}\to\Omega_{c}^{(*)0}M$ with $M=(\pi^{-},\rho^{-})$ proceed through the quark-level $b\to c\bar{u}d$ transition. The according effective Hamiltonian is given by Buchalla:1995vs
${\cal H}_{eff}=\frac{G_{F}}{\sqrt{2}}V_{cb}V_{ud}^{\ast}(c_{1}O_{1}+c_{2}O_{2})\,$, (1)
where $G_{F}$ is the Fermi constant, $c_{1,2}$ are the Wilson coefficients, and $V_{cb}(V_{ud}^{*})$ the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements. Explicitly, the current-current operators $O_{1,2}$ read
$O_{1}$ $=$ $\bar{\psi}_{\bar{u}_{\beta}}\gamma_{\mu}(1-\gamma_{5})\psi_{d_{\beta}}\bar{\psi}_{\bar{c}_{\alpha}}\gamma^{\mu}(1-\gamma_{5})\psi_{b_{\alpha}}\,$,
$O_{2}$ $=$ $\bar{\psi}_{\bar{u}_{\alpha}}\gamma_{\mu}(1-\gamma_{5})\psi_{d_{\beta}}\bar{\psi}_{\bar{c}_{\beta}}\gamma^{\mu}(1-\gamma_{5})\psi_{b_{\alpha}},$ (2)
where $\psi_{j_{\delta}}$ represent the $j$-quark field $[j=(u,d,c,b)]$ and $\delta=(\alpha,\beta)$ denotes a color quantum number. The amplitude of $\Omega_{b}^{-}\to\Omega_{c}^{(*)0}M$ through ${\cal H}_{eff}$ in Eq. (1) is derived as Hsiao:2015txa ; Hsiao:2015cda ; Hsiao:2021mlp
${\cal M}=\frac{G_{F}}{\sqrt{2}}V_{cb}V_{ud}^{\ast}a_{1}\langle M\Omega_{c}^{(*)0}|\bar{\psi}_{\bar{u}_{\beta}}\gamma_{\mu}(1-\gamma_{5})\psi_{d_{\beta}}\bar{\psi}_{\bar{c}_{\alpha}}\gamma^{\mu}(1-\gamma_{5})\psi_{b_{\alpha}}|\Omega_{b}^{-}\rangle\,$, (3)
with $a_{1}=c_{1}^{eff}+c_{2}^{eff}/N_{c}^{eff}$, where $c_{1,2}^{eff}$ denote the effective Wilson coefficient that receive the quark-rescattering effects, and $N_{c}^{eff}$ denote the effective color number. By shifting $N_{c}^{eff}$ from $N_{c}=3$, the non-factorizable QCD loop effects can be estimated ali ; Hsiao:2014mua . In Eq. (3), the two currents in $O_{1}$ are split into the two matrix elements, and $O_{2}\simeq O_{1}/N_{c}^{(eff)}$ as a result of the factorization approximation has been combined as part of the $O_{1}$ contribution.
Figure 1: Feynman diagram for the processes of $\Omega_{b}\rightarrow\Omega_{c}^{(\ast)}{M}$.
The amplitude in Eq. (3) can be further derived, where $O_{1}$ being split into the two matrix elements is presented as $O_{1}\simeq O_{\rm w}^{PC}+O_{\rm w}^{PV}$ in the non-relativistic framework. The parity-conserving and parity-violating operators of the weak transition, denoted by $O_{\rm w}^{PV}$ and $O_{\rm w}^{PV}$, respectively, are given by
$O_{\rm w}^{PC}$ $=$ $\delta^{3}(\textbf{p}_{3}-\textbf{p}_{3}^{\prime}-\textbf{p}_{4}-\textbf{p}_{5})/(2\pi)^{3}\hat{O}_{f}\hat{O}_{c}$
$\times\bigg\{$ $\mbox{$\sigma$}_{4}\cdot\bigg[\left(\frac{\textbf{p}_{5}}{2m_{5}}+\frac{\textbf{p}_{4}}{2m_{4}}\right)-\left(\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}+\frac{\textbf{p}_{3}}{2m_{3}}\right)+i\mbox{$\sigma$}_{3}\times\left(\frac{\textbf{p}_{3}}{2m_{3}}-\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}\right)\bigg]$
$+$ $\mbox{$\sigma$}_{3}\cdot\bigg[\left(\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}+\frac{\textbf{p}_{3}}{2m_{3}}\right)-\left(\frac{\textbf{p}_{5}}{2m_{5}}+\frac{\textbf{p}_{4}}{2m_{4}}\right)+i\mbox{$\sigma$}_{4}\times\left(\frac{\textbf{p}_{4}}{2m_{4}}-\frac{\textbf{p}_{5}}{2m_{5}}\right)\bigg]\bigg\}\,$,
$O_{\rm w}^{PV}$ $=$ $\delta^{3}(\textbf{p}_{3}-\textbf{p}_{3}^{\prime}-\textbf{p}_{4}-\textbf{p}_{5})/(2\pi)^{3}\hat{O}_{f}\hat{O}_{c}(\mbox{$\sigma$}_{3}\cdot\mbox{$\sigma$}_{4}-1)\,$, (4)
where $\textbf{p}_{j}$, $m_{j}$, and $\mbox{$\sigma$}_{j}$ as the momentum, mass, and spin of the $j$-quark, respectively, are assigned as in Fig. 1. In Eq. (II), $\hat{O}_{f}=b_{5}^{\dagger}(d)b_{4}^{\dagger}(\bar{u})b_{3}^{\dagger}(c)b_{3}(b)$ transforms $b$ to $c$ quark for the $\Omega_{b}^{-}\to\Omega_{c}^{(*)}$ transition, and creates the $\bar{u}d$ pair for the final-state meson. Moreover, $\hat{O}_{c}=\delta_{c_{4},c_{5}}\delta_{c_{3}^{\prime},c_{3}}$ is the color operator, where $\delta_{c_{m},c_{n}}$ corresponds to the $W$-boson emission, which is color-singlet.
In the constituent quark model, the meson and baryon wave functions are established as
$|M(\mathbf{P}_{M},J,J_{z})\rangle$ $=$ $\sum_{S_{z},L_{z}}\langle L,L_{z};S,S_{z}|J,J_{z}\rangle\int d\mathbf{p}_{4}d\mathbf{p}_{5}\delta^{3}(\mathbf{p}_{4}+\mathbf{p}_{5}-\mathbf{P}_{M})$
$\times$ $\Psi_{M}(\mathbf{p}_{4},\mathbf{p}_{5})|q_{4}(\mathbf{p}_{4})q_{5}(\mathbf{p}_{5})\rangle\,$,
$|{\bf B}(\mathbf{P}_{\bf B},J,J_{z})\rangle$ $=$ $\sum_{S_{z},L_{z}}\langle L,L_{z};S,S_{z}|J,J_{z}\rangle\int d\mathbf{p}_{1}d\mathbf{p}_{2}d\mathbf{p}_{3}\delta^{3}(\mathbf{p}_{1}+\mathbf{p}_{2}+\mathbf{p}_{3}-\mathbf{P}_{\bf B})\,$ (5)
$\times$ $\Psi_{\bf B}(\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3})|q_{1}(\mathbf{p}_{1})q_{2}(\mathbf{p}_{2})q_{3}(\mathbf{p}_{3})\rangle\,$,
which correspond to the kets $|M\rangle$ and $|\Omega_{b}\rangle$ in Eq. (3), respectively. By adding a prime notation in $\mathbf{p}_{3}$, $|{\bf B}(\mathbf{P}_{\bf B},J,J_{z})\rangle$ correspond to $\Omega_{c}^{(*)0}$. The meson wave function $\Psi_{M}(\mathbf{p}_{4},\mathbf{p}_{5})$ in the momentum space is given by
$\Psi_{{M}}(\textbf{p}_{4},\textbf{p}_{5})=\zeta_{{M}}\varphi_{{M}}\chi^{S}_{S_{z}}\psi_{M}(\textbf{p}_{4},\textbf{p}_{5})\,$, (6)
where $\zeta_{M}$, $\varphi_{M}$, and $\chi^{S}_{S_{z}}$ stand for the color, flavor, and spin wave functions, respectively. For the final-state mesons $\pi^{-}$ and $\rho^{-}$, we present $\zeta_{\pi^{-},\rho^{-}}=(R\bar{R}+G\bar{G}+B\bar{B})/\sqrt{3}$ and $\varphi_{\pi^{-},\rho^{-}}=\bar{u}d$, while $\chi^{0}_{0}=(\uparrow\downarrow-\downarrow\uparrow)/\sqrt{2}$ and $\chi^{1}_{1,0,-1}=[\uparrow\uparrow,(\uparrow\downarrow+\downarrow\uparrow)/\sqrt{2},\downarrow\downarrow]$ are the $\pi^{-}$ and $\rho^{-}$ spin wave functions, respectively. As the spatial wave function, $\psi_{M}(\textbf{p}_{4},\textbf{p}_{5})$ is written as
$\psi_{\pi(\rho)}(\textbf{p}_{4},\textbf{p}_{5})=\frac{2^{3/4}}{\pi^{3/4}\beta_{\pi(\rho)}^{3/2}}\mathrm{exp}\left[-\frac{(\textbf{p}_{4}-\textbf{p}_{5})^{2}}{4\beta_{\pi(\rho)}^{2}}\right]\,$, (7)
where the coefficient $\beta_{\pi(\rho)}$ controls the Gaussian distribution.
The baryon wave function $\Psi_{\bf B}(\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3})$ can be more complicated, given by
$\Psi_{\bf B}(\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3})=\zeta_{\bf B}\varphi_{\bf B}\chi^{S}_{S_{z}}\Psi_{NLM_{L}}(\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3})\,$, (8)
where $\zeta_{\bf B}$, $\varphi_{\bf B}$, $\chi^{S}_{S_{z}}$, and $\Psi_{NLM_{L}}$ represent the color, flavor, spin, and spatial wave functions, respectively. For the $\Omega_{b(c)}$ and $\Omega_{c}^{(*)}$ baryons, the corresponding color and flavor wave functions are $\zeta_{\Omega_{b}}=\zeta_{\Omega_{c}^{(*)}}=(RGB-RBG+GBR-GRB+BRG-BGR)/\sqrt{6}$ and $\varphi_{\Omega_{b},\Omega_{c}^{(*)}}=(bss,css)$.
The spin and spatial wave functions of a baryon are not independent, since both are determined by the internal three-quark dynamics. To illustrate this relation, one may adopt the Jacobi-momentum framework, in which the three-momenta $\mathbf{p}_{1,2,3}$ of the constituent quarks $q_{1,2,3}$ are reorganized as
$\mathbf{p}_{\rho}=\frac{\sqrt{2}}{2}(\textbf{p}_{1}-\textbf{p}_{2})\,,\;\mathbf{p}_{\lambda}=\frac{\sqrt{6}}{2}\frac{m_{3}(\textbf{p}_{1}+\textbf{p}_{2})-(m_{1}+m_{2})\textbf{p}_{3}}{m_{1}+m_{2}+m_{3}}\,.$ (9)
Here, $\mathbf{p}_{\rho}$ describes the relative motion between $q_{1}$ and $q_{2}$, while $\mathbf{p}_{\lambda}$ characterizes the motion of the third quark $q_{3}$ with respect to the $q_{1}$-$q_{2}$ subsystem. Accordingly, the $q_{1}q_{2}$ pair forms a diquark associated with the $\rho$-mode oscillator, whereas the motion between $q_{3}$ and the diquark corresponds to the $\lambda$-mode oscillator. The baryon wave function, originally written as $\chi^{S}_{S_{z},s_{\rho}}\Psi_{NLM_{L}}(\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3})$, can therefore be expressed in the Jacobi basis as $\chi^{S}_{S_{z},s_{\rho}}\Psi_{NLM_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})$, where the $\rho$ label indicates that the spin structure of the $q_{1}q_{2}$ subsystem is correlated with the corresponding $\rho$-mode spatial excitation.
The $\Omega_{b}$ and the ground-state $\Omega_{c}$ baryons are both spin-1/2 fermions. In these systems, the two strange quarks form a symmetric ssss-diquark with spin $s_{\rho}=1$, leading to the total spin $S=s_{\rho}-s_{c}=1/2$, where $s_{c}=1/2$ denotes the charm quark spin. Accordingly, their spin wave functions can be written as $\chi^{1/2}_{1/2,1}=-1/\sqrt{6}(\uparrow\downarrow\uparrow+\downarrow\uparrow\uparrow-2\uparrow\uparrow\downarrow)$ and $\chi^{1/2}_{-1/2,1}=1/\sqrt{6}(\uparrow\downarrow\downarrow+\downarrow\uparrow\downarrow-2\downarrow\downarrow\uparrow)$. As another example, $\Omega_{c}(2770)$ also contains a spin-1 ssss diquark ($s_{\rho}=1$) but carries total spin $S=s_{\rho}+s_{c}=3/2$. Its spin wave-function components are given by $\chi^{3/2}_{3/2,1}=\uparrow\uparrow\uparrow$, $\chi^{3/2}_{1/2,1}=(\uparrow\uparrow\downarrow+\uparrow\downarrow\uparrow+\downarrow\uparrow\uparrow)/\sqrt{3}$, $\chi^{3/2}_{-1/2,1}=1/\sqrt{3}(\uparrow\downarrow\downarrow+\downarrow\uparrow\downarrow+\downarrow\downarrow\uparrow)$, and $\chi^{3/2}_{-3/2,1}=\downarrow\downarrow\downarrow$.
Table 1: The spin-space wave functions of the $\Omega_{c}$ baryons up to the $1D$-wave states, along with the assignment of the quantum numbers. The configurations are labeled in the form $|N^{2S+1}LJ^{P}\rangle$.
$\Omega_{c}$ baryon states NN $n_{\rho}$ $n_{\lambda}$ $L$ $S$ $s_{\rho}$ $s_{c}$ $l_{\rho}$ $l_{\lambda}$ $J^{P}$ Wave function
$|\Omega_{c}~1^{4}S\frac{3}{2}^{+}\rangle$ 11 0 0 0 $\frac{3}{2}$ 11 $\frac{1}{2}$ 0 0 $\frac{3}{2}^{+}$ $\chi^{3/2}_{S_{z},1}\psi_{000}(\textbf{p}_{\rho})\psi_{000}(\textbf{p}_{\lambda})$
$|\Omega_{c}~1^{2}S\frac{1}{2}^{+}\rangle$ 11 0 0 0 $\frac{1}{2}$ 11 $\frac{1}{2}$ 0 0 $\frac{1}{2}^{+}$ $\chi^{1/2}_{S_{z},1}\psi_{000}(\textbf{p}_{\rho})\psi_{000}(\textbf{p}_{\lambda})$
$|\Omega_{c}~2^{4}S_{\lambda}\frac{3}{2}^{+}\rangle$ 22 0 1 0 $\frac{3}{2}$ 11 $\frac{1}{2}$ 0 0 $\frac{3}{2}^{+}$ $\chi^{3/2}_{S_{z},1}\psi_{000}(\textbf{p}_{\rho})\psi_{100}(\textbf{p}_{\lambda})$
$|\Omega_{c}~2^{2}S_{\lambda}\frac{1}{2}^{+}\rangle$ 22 0 1 0 $\frac{1}{2}$ 11 $\frac{1}{2}$ 0 0 $\frac{1}{2}^{+}$ $\chi^{1/2}_{S_{z},1}\psi_{000}(\textbf{p}_{\rho})\psi_{100}(\textbf{p}_{\lambda})$
$|\Omega_{c}~1^{4}P_{\lambda}J^{-}\rangle$ 11 0 0 1 $\frac{3}{2}$ 11 $\frac{1}{2}$ 0 11 $\frac{5}{2}^{-},\frac{3}{2}^{-},\frac{1}{2}^{-}$ $\chi^{3/2}_{S_{z},1}\psi_{000}(\textbf{p}_{\rho})\psi_{01m_{\lambda}}(\textbf{p}_{\lambda})$
$|\Omega_{c}~1^{2}P_{\lambda}J^{-}\rangle$ 11 0 0 1 $\frac{1}{2}$ 11 $\frac{1}{2}$ 0 11 $\frac{3}{2}^{-},\frac{1}{2}^{-}$ $\chi^{1/2}_{S_{z},1}\psi_{000}(\textbf{p}_{\rho})\psi_{01m_{\lambda}}(\textbf{p}_{\lambda})$
$|\Omega_{c}~1^{4}D_{\lambda}J^{+}\rangle$ 11 0 0 2 $\frac{3}{2}$ 11 $\frac{1}{2}$ 0 22 $\frac{7}{2}^{+},\frac{5}{2}^{+},\frac{3}{2}^{+},\frac{1}{2}^{+}$ $\chi^{3/2}_{S_{z},1}\psi_{000}(\textbf{p}_{\rho})\psi_{02m_{\lambda}}(\textbf{p}_{\lambda})$
$|\Omega_{c}~1^{2}D_{\lambda}J^{+}\rangle$ 11 0 0 2 $\frac{1}{2}$ 11 $\frac{1}{2}$ 0 22 $\frac{5}{2}^{+},\frac{3}{2}^{+}$ $\chi^{1/2}_{S_{z},1}\psi_{000}(\textbf{p}_{\rho})\psi_{02m_{\lambda}}(\textbf{p}_{\lambda})$
In the $\rho$-$\lambda$ oscillation model, the spatial wave function of a three-quark system is expressed as Weng:2024roa
$\Psi_{NLL_{z}}(\textbf{p}_{\rho},\textbf{p}_{\lambda})=\sum_{m_{\rho},m_{\lambda}}\langle l_{\rho},m_{\rho};l_{\lambda},m_{\lambda}|L,L_{z}\rangle\psi_{n_{\rho}l_{\rho}m_{\rho}}(\textbf{p}_{\rho})\psi_{n_{\lambda}l_{\lambda}m_{\lambda}}(\textbf{p}_{\lambda})\,$, (10)
where the principal quantum number is defined as $N=n_{\rho}+n_{\lambda}+1$, the total angular momentum satisfies $L=l_{\rho}+l_{\lambda},...,|l_{\rho}-l_{\lambda}|$, and $L_{z}=m_{\rho}+m_{\lambda}$. Here, $(n_{i},l_{i},m_{i}$) with $i=\rho$ or $\lambda$ denote the radial, orbital, and magnetic quantum numbers, respectively, in the $\rho$ or $\lambda$ oscillation system. Assuming the simple harmonic oscillator wave functions, $\psi_{n_{\rho}l_{\rho}m_{\rho}}(\textbf{p}_{\rho})$ and $\psi_{n_{\lambda}l_{\lambda}m_{\lambda}}(\textbf{p}_{\lambda})$ in Eq. (10) share the same functional form,
$\psi_{nlm}(\textbf{p})=(i)^{l}(-1)^{n}\left[\frac{2n!}{(n+l+1/2)!}\right]^{1/2}\frac{1}{\alpha^{l+3/2}}\mathrm{exp}\left(-\frac{\textbf{p}^{2}}{2\alpha^{2}}\right)L_{n}^{l+1/2}(\textbf{p}^{2}/\alpha^{2})\mathcal{Y}_{lm}(\textbf{p})\,$, (11)
[…]
With these ingredients, the spin-space wave functions of the $\Omega_{c}$ baryons can be constructed, as summarized in Table 1. In particular, the notation $|N^{2S+1}LJ^{P}\rangle$ specifies the quark configuration of the $\Omega_{c}$ baryon in the $\rho$-$\lambda$ scheme, where the orbital angular momentum $L=0,1,2$ corresponds to the $S$-, $P$-, and $D$-wave states, respectively. The total spin is given by $S=s_{\rho}\pm s_{c}$, and the total angular momentum is obtained from $J=L+S$. The subscript $\lambda$ in $S_{\lambda}$ or $P_{\lambda}$ ($D_{\lambda}$) indicates that the excitation arises from the $\lambda$ mode, corresponding to $N=n_{\lambda}+1$ (radial excitation) or $L=l_{\lambda}$ (orbital excitation) in the $\lambda$ oscillator. These constructions lead to the spin-space wave functions (see the last column of Table 1).
In the constituent quark model, the transition amplitude $\langle\Omega_{c}^{(*)0}M|O_{1}|\Omega_{b}^{-}\rangle$ with the insertion of $O_{\rm w}^{PC,PV}$ in Eq. (II) and the wave functions in Eq. (II) is evaluated as
$\mathcal{M}_{J_{f},J_{f}^{z};J_{{M}},J_{{M}}^{z}}^{J_{i},J_{i}^{z}}(\Omega_{b}^{-}\to\Omega_{c}^{(*)}{M})=\mathcal{M}_{J_{f},J_{f}^{z};J_{{M}},J_{{M}}^{z}}^{J_{i},J_{i}^{z},PC}+\mathcal{M}_{J_{f},J_{f}^{z};J_{{M}},J_{{M}}^{z}}^{J_{i},J_{i}^{z},PV}\,$,
$\mathcal{M}_{J_{f},J_{f}^{z};J_{{M}},J_{{M}}^{z}}^{J_{i},J_{i}^{z},PC(PV)}=\langle\Omega_{c}^{(*)}(\textbf{P}_{f},J_{f},J_{f}^{z}){M}(\textbf{q},J_{{M}},J_{{M}}^{z})|O_{\rm w}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i},J_{i},J_{i}^{z})\rangle\,$, (12)
where $J_{i,f,M}^{(z)}$ denote the total (third) angular momenta of the initial-state baryon, final baryon, and final meson, respectively, and $\textbf{q}=\textbf{P}_{i}-\textbf{P}_{f}$ is the momentum transfer.
As an explicit example in our calculation, we consider the decay $\Omega_{b}^{-}\rightarrow\Omega_{c}^{0}\pi^{-}$. The corresponding angular-momentum assignments are $(J_{f},J_{{M}},J_{i})=(1/2,0,1/2)$ and $(J_{f}^{z},J_{{M}}^{z},J_{i}^{z})=(-1/2,0,-1/2)$. Using Eq. (II), the parity-conserving (PC) and parity-violating (PV) amplitudes are obtained as
$\mathcal{M}_{1/2,-1/2,0,0}^{1/2,-1/2,PC}$ $=$ $-\frac{\sqrt{2}}{\sqrt{3}\pi^{9/4}}a_{1}G_{F}V_{cb}V_{ud}^{\ast}\frac{T_{\alpha}\beta_{\pi}^{9/2}}{(\alpha_{\lambda_{f}}^{2}+\alpha_{\lambda_{i}}^{2})(2m_{2}+m_{3}^{\prime})}qF(q^{2})$
$\times\bigg[\alpha_{\lambda_{i}}^{2}\left(1-\frac{2m_{2}}{m_{3}}+\frac{m_{2}}{m_{4}}+\frac{m_{2}}{m_{5}}+\frac{m_{3}^{\prime}}{2m_{4}}+\frac{m_{3}^{\prime}}{2m_{5}}\right)$
$+\alpha_{\lambda_{f}}^{2}\left(1+\frac{{m_{2}}}{m_{4}}+\frac{m_{2}}{m_{5}}+\frac{2m_{2}}{m_{3}^{\prime}}+\frac{m_{3}^{\prime}}{2m_{4}}+\frac{m_{3}^{\prime}}{2m_{5}}\right)\bigg]\,$,
$\mathcal{M}_{1/2,-1/2,0,0}^{1/2,-1/2,PV}$ $=$ $-\frac{2\sqrt{6}}{\pi^{9/4}}a_{1}G_{F}V_{cb}V_{ud}^{\ast}\beta_{\pi}^{9/2}T_{\alpha}F(q^{2})\,$, (13)
where the oscillator parameters are defined by $\alpha_{\lambda_{i}}=[3m_{b}/(2m_{s}+m_{b})]^{1/4}\alpha_{\rho_{i}}$ and $\alpha_{\lambda_{f}}=[3m_{c}/(2m_{s}+m_{c})]^{1/4}\alpha_{\rho_{f}}$. The quantities $T_{\alpha}$ and $F(q^{2})$ are given by
$T_{\alpha}=\left[\frac{\alpha_{\lambda_{f}}\alpha_{\lambda_{i}}\alpha_{\rho_{f}}\alpha_{\rho_{i}}}{\beta_{\pi}^{2}\left(\alpha_{\lambda_{f}}^{2}+\alpha_{\lambda_{i}}^{2}\right)\left(\alpha_{\rho_{f}}^{2}+\alpha_{\rho_{i}}^{2}\right)}\right]^{3/2},F(q^{2})=\exp\left[\frac{-3m_{2}^{2}q^{2}}{\left(\alpha_{\lambda_{f}}^{2}+\alpha_{\lambda_{i}}^{2}\right)(2m_{2}+m_{3}^{\prime})^{2}}\right]\,.$ (14)
Using the equation
$\Gamma$ $=$ $8\pi^{2}\frac{|\textbf{q}|E_{{M}}E_{f}}{M_{i}}\frac{1}{2J_{i}+1}\sum_{J_{f}^{z},J_{i}^{z}}|\mathcal{M}_{J_{f},J_{f}^{z};J_{{M}},J_{{M}}^{z}}^{J_{i},J_{i}^{z}}(\Omega_{b}\to{M}\Omega_{c}^{(*)})|^{2}\,$, (15)
the helicity amplitudes are converted into decay widths for the numerical analysis. Here, $\mathbf{q}$ denotes the three-momentum of the final-state particles in the $\Omega_{b}$ rest frame, and the spin projections $J_{i}^{z}$ and $J_{f}^{z}$ are summed over in the squared amplitudes, with the factor $1/(2J_{i}+1)$ accounting for the average over the initial-state polarization. As a result, the PC and PV amplitudes for the decay $\Omega_{b}^{-}\rightarrow\Omega_{c}^{0}\pi^{-}$, as constructed in Eq. (II), are directly substituted into Eq. (15) to obtain the corresponding decay width. The same procedure is applied to compute all other decay widths considered in this work.
III Numerical Analysis
In the numerical analysis, we adopt the CKM matrix elements $(V_{cb},V_{ud})=(0.042,0.978)$ from the PDG pdg and the quark masses $(m_{s},m_{c},m_{b})=(0.55,1.45,4.80)$ GeV from Refs. Ni:2023lvx ; Wang:2025pcs , together with the parameter $a_{1}=1.1$ taken from Ref. Gutsche:2018utw ; Wang:2024mjw . The harmonic oscillator parameters $\alpha_{\rho_{i}(\rho_{f})}$ and $\beta_{\pi(\rho)}$ are taken from Ref. Wang:2017kfr and Refs. Zhong:2008kd ; Wang:2025pcs , respectively, with values
$\alpha_{\rho_{i}}=\alpha_{\rho_{f}}=440~\text{MeV}\,$,\;
$\beta_{\pi}=\beta_{\rho}=400~\text{MeV}\,.$ (16)
Using the parameters in Eq. (III) as theoretical inputs, together with $O_{\rm w}^{PC,PV}$ in Eq. (II), the wave functions in Eqs. (II, 10, 11), and the amplitudes in Eq. (II), we evaluate the branching fractions ${\cal B}(\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-},\Omega_{c}^{0}\rho^{-})$, where $\Omega_{c}^{0}$ is the lowest-lying $1S$-wave $\Omega_{c}$ state. The numerical results are presented in Table 2, together with a comparison with other model predictions. We further predict the branching fractions for $\Omega_{b}\to\Omega_{c}^{*}M$, including final-state $\Omega_{c}$ baryons in the $1S$-, $1P$-, $2S$-, and $1D$-wave configurations.
In the LLSS coupling scheme, the $\Omega_{c}$ wave states are labeled by $|LJ^{P}\rangle$, as listed in Table 1. In this framework, the total orbital angular momentum $L$ is obtained by coupling the $\rho$- and $\lambda$-mode orbital momenta, $L={\ell}_{\rho}+{\ell}_{\lambda}$, while the total spin is given by $S=s_{\rho}+s_{c}$. The total angular momentum of the charmed baryon $\Omega_{c}^{(*)}$ is then obtained as $J=L+S$. It is worth noting that, in the heavy-quark limit, the charm-quark spin $s_{c}$ decouples from the light degrees of freedom. This motivates the use of a more suitable framework, in which the $\Omega_{c}$ wave states are labeled by $|J^{P},j_{\ell}\rangle$ in the jjj-j coupling scheme Wang:2017hej ; Wang:2018fjm ; Xiao:2020oif . In this notation, $j_{\ell}=L+s_{\rho}$ denotes the total angular momentum of the light-quark system. The total angular momentum of the baryon is then obtained by coupling $j_{\ell}$ with the charm-quark spin, $J=j_{\ell}+s_{c}$, for the $\Omega_{c}^{(*)}$ states considered in this work.
The two coupling schemes are related by a unitary transformation Roberts:2007ni ,
$|J^{P},j_{\ell}\rangle=\sum_{S}(-1)^{L+s_{\rho}+s_{c}+J}\sqrt{(2S+1)(2j_{\ell}+1)}\begin{Bmatrix}s_{\rho}&s_{c}&S\\ L&J&j_{\ell}\end{Bmatrix}|LJ^{P}\rangle\,$, (17)
where the bracket denotes the Wigner $6j$ symbol describing the recoupling of angular momenta. In the jjjj coupling scheme, the charmed baryons $\Omega_{c}^{(*)}$ considered in this work are labeled by $|J^{P},j_{\ell}\rangle$, as summarized in Table 3. Moreover, the subscript $\lambda$ in $|J^{P},j_{\ell}\rangle_{\lambda}$ for the $1P$-, $2S$-, and $1D$-wave states indicates that the excitation occurs in the $\lambda$ mode, while the $\rho$ mode remains in its ground state.
Table 2: Branching fractions of $\Omega_{b}^{-}\rightarrow\Omega_{c}^{0}\pi^{-}$ and $\Omega_{b}^{-}\rightarrow\Omega_{c}^{0}\rho^{-}$, in comparison with previous works.
$10^{3}\mathcal{B}(\Omega_{b}^{-}\rightarrow\Omega_{c}^{0}\pi^{-})$ $10^{3}\mathcal{B}(\Omega_{b}^{-}\rightarrow\Omega_{c}^{0}\rho^{-})$
CQM (this work) 1.2 6.3
LFQM Chua:2019yqh $1.1^{+0.85}_{-0.55}$ $3.07^{+2.41}_{-1.53}$
LFQM Zhao:2018zcb 4.0 10.8
LFQM Li:2021kfb 2.82 7.92
QCDF Neishabouri:2026mjn 1.809 5.047
RQM Ivanov:1997ra ; Ivanov:1997hi 5.81 $\cdots$
RQM Patel:2025gbw 2.54 7.43
CCQM Gutsche:2018utw 1.88 5.4
NRQM Cheng:1996cs 4.92 17.9
IV Discussions and Conclusion
Experimentally, the decay $\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-}$ has been observed through its subsequent decay $\Omega_{c}^{0}\to pK^{-}K^{-}\pi^{+}$ LHCb:2016coe . However, the corresponding branching fraction has not yet been reported. Consequently, experimental data are currently insufficient to test the validity of theoretical models for processes involving the ground-state $\Omega_{c}^{0}$. On the other hand, as shown in Table 2, our predicted branching fractions ${\cal B}(\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-},\Omega_{c}^{0}\rho^{-})=(1.2,6.3)\times 10^{-3}$ are consistent, at least at the order-of-magnitude level, with the existing calculations in the light-front quark model (LFQM) Chua:2019yqh ; Zhao:2018zcb ; Li:2021kfb , the relativistic quark model (RQM) Ivanov:1997ra ; Ivanov:1997hi ; Patel:2025gbw , the QCD factorization approach Patel:2025gbw , and the covariant confined quark model (CCQM) Gutsche:2018utw . This agreement suggests that the constituent quark model employed in this work provides a reasonable framework comparable to other theoretical approaches. Based on this consistency, we further present our predictions for $\Omega_{b}^{-}\to\Omega_{c}^{*0}\pi^{-},\Omega_{c}^{*0}\rho^{-}$.
In the $1S$-wave classification, there exists an excited state, $\Omega_{c}(2770)^{0}$, with spin $3/2$. Its existence has been experimentally established, and it is found to decay dominantly through the radiative channel. Using the branching fractions ${\cal B}(\Omega_{b}^{-}\to\Omega_{c}(2770)^{0}\pi^{-})$ and ${\cal B}(\Omega_{b}^{-}\to\Omega_{c}(2770)^{0}\rho^{-})$ predicted in Table 3, together with ${\cal B}(\Omega_{c}(2770)^{0}\to\Omega_{c}^{0}\gamma)\simeq 100\%$ from the PDG pdg , we estimate the cascade radiative branching fractions
${\cal B}(\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-}\gamma,\Omega_{c}^{0}\rho^{-}\gamma)\simeq(5.2,1.4)\times 10^{-3}\,.$ (18)
In addition, radiative processes $\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-}\gamma$ and $\Omega_{b}^{-}\to\Omega_{c}^{0}\rho^{-}\gamma$ can also arise from photon emission via inner bremsstrahlung (IB) from the charged external legs in the non-radiative decays $\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-}$ and $\Omega_{b}^{-}\to\Omega_{c}^{0}\rho^{-}$. Since the IB contribution is infrared divergent, we implicitly assume a photon-energy threshold $E_{\gamma}>E_{0}$ in the $\Omega_{b}$ rest frame. A rough estimate gives ${\cal B}_{\rm IB}\sim\delta_{\rm IB}\,{\cal B}$ with $\delta_{\rm IB}={\cal O}\!\left[(\alpha_{\rm em}/\pi)\ln(E_{\rm hard}/E_{0})\right]\sim 10^{-2}$ Geng:2004jk for typical experimental cuts, leading to ${\cal B}_{\rm IB}(\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-}\gamma,\Omega_{c}^{0}\rho^{-}\gamma)\sim 10^{-5}$. By comparison, the cascade contributions in Eq. (18) are significantly larger than the IB background, indicating that these radiative $\Omega_{b}$ decay modes may provide promising channels for future experimental measurements.
Table 3: Branching fractions of $\Omega_{b}\to\Omega_{c}^{(*)}{M}$ decay processes. The final-state $\Omega_{c}^{(*)}$ are classified, from top to button, as the $1S$-, $1P$-, $2S$-, and $1D$-wave states. The corresponding masses are denoted by $M_{f}$, while $|J^{P},j_{\ell}\rangle$ specifies the wave state of the $\Omega_{c}^{(*)}$ baryon. We define ${\cal B}_{M}={\cal B}(\Omega_{b}\to\Omega_{c}^{(*)}M)$ to present our predicted branching fractions.
wave $\Omega_{c}$ baryons $|J^{P},j_{\ell}\rangle$ NN $n_{\lambda}$ $n_{\rho}$ $l_{\lambda}$ $l_{\rho}$ $L$ $s_{\rho}$ $j_{\ell}$ $J^{P}$ $M_{f}$ (MeV) $10^{3}\mathcal{B}_{\pi}$ $10^{3}\mathcal{B}_{\rho}$
1S $\Omega_{c}^{0}$ $|{J^{P}=\frac{1}{2}^{+},1}\rangle$} 1 0 0 0 0 0 1 1 $\frac{1}{2}^{+}$ 2695 1.2 6.3
$\Omega_{c}(2770)$ $|{J^{P}=\frac{3}{2}^{+},1}\rangle$} 1 0 0 0 0 0 1 1 $\frac{3}{2}^{+}$ 2770 5.2 1.4
1P $\Omega_{c}^{\prime}(3000)$ ${|{J^{P}=\frac{1}{2}^{-},0}\rangle}_{\lambda}$ 1 0 0 1 0 1 1 0 $\frac{1}{2}^{-}$ 3000 2.2 2.0
$\Omega_{c}(3050)$ ${|{J^{P}=\frac{3}{2}^{-},1}\rangle}_{\lambda}$ 1 0 0 1 0 1 1 1 $\frac{3}{2}^{-}$ 3055 1.9 2.3
$\Omega_{c}(3000)$ ${|{J^{P}=\frac{1}{2}^{-},1}\rangle}_{\lambda}$ 1 0 0 1 0 1 1 1 $\frac{1}{2}^{-}$ 3000 4.4 4.0
$\Omega_{c}(3065)$ ${|{J^{P}=\frac{3}{2}^{-},2}\rangle}_{\lambda}$ 1 0 0 1 0 1 1 2 $\frac{3}{2}^{-}$ 3066 1.3 9.8
$\Omega_{c}(3090)$ ${|{J^{P}=\frac{5}{2}^{-},2}\rangle}_{\lambda}$ 1 0 0 1 0 1 1 2 $\frac{5}{2}^{-}$ 3090 8.1 1.5
2S *** ${|{J^{P}=\frac{1}{2}^{+},1}\rangle}_{\lambda}$ 2 1 0 0 0 0 1 1 $\frac{1}{2}^{+}$ 3088 1.5 7.3
$\Omega_{c}(3185)$ ${|{J^{P}=\frac{3}{2}^{+},1}\rangle}_{\lambda}$ 2 1 0 0 0 0 1 1 $\frac{3}{2}^{+}$ 3177 6.0 1.4
1D *** ${|{J^{P}=\frac{3}{2}^{+},1}\rangle}_{\lambda}$ 1 0 0 2 0 2 1 1 $\frac{3}{2}^{+}$ 3296 Ebert:2011kk 0.6 2.0
*** ${|{J^{P}=\frac{5}{2}^{+},2}\rangle}_{\lambda}$ 1 0 0 2 0 2 1 2 $\frac{5}{2}^{+}$ 3305 Ebert:2011kk 1.5 1.6
*** ${|{J^{P}=\frac{1}{2}^{+},1}\rangle}_{\lambda}$ 1 0 0 2 0 2 1 1 $\frac{1}{2}^{+}$ 3264 Ebert:2011kk 7.5 0.3
*** ${|{J^{P}=\frac{3}{2}^{+},2}\rangle}_{\lambda}$ 1 0 0 2 0 2 1 2 $\frac{3}{2}^{+}$ 3280 Ebert:2011kk 3.0 2.6
$\Omega_{c}(3327?)$ ${|{J^{P}=\frac{5}{2}^{+},3}\rangle}_{\lambda}$ 1 0 0 2 0 2 1 3 $\frac{5}{2}^{+}$ 3296 Ebert:2011kk 2.0 5.7
$\Omega_{c}(3327?)$ ${|{J^{P}=\frac{7}{2}^{+},3}\rangle}_{\lambda}$ 1 0 0 2 0 2 1 3 $\frac{7}{2}^{+}$ 3306 Ebert:2011kk 4.6 0.8
Four of the five observed $\Omega_{c}^{**}$ states can be classified as the $1P$-wave excitations. In particular, $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$, and $\Omega_{c}(3090)$ are observed as resonant peaks in the $\Xi_{c}^{+}K^{-}$ invariant mass spectrum from the cascade process $\Omega_{b}^{-}\to\Omega_{c}^{**}\pi^{-}$ followed by $\Omega_{c}^{**}\to\Xi_{c}^{+}K^{-}$ LHCb:2021ptx . By contrast, $\Omega_{c}(3120)$ has not yet been experimentally confirmed. Motivated by its non-observation, $\Omega_{c}(3120)$ may be interpreted as a state dominated by a $\rho$-mode excitation Zhong:2025oti ; Weng:2024roa . Since the initial $\Omega_{b}$ state is characterized by $n_{\rho}=\ell_{\rho}=0$ and therefore cannot efficiently undergo a weak transition to $\rho$-mode excitations, it is reasonable that the $\Omega_{b}$ decay into $\Omega_{c}(3120)$ is forbidden or highly suppressed.
Despite the experimental observation of the processes $\Omega_{b}^{-}\to\Omega_{c}^{**0}\pi^{-}$, their branching fractions have not yet been reported. We therefore present our predictions, together with those for ${\cal B}(\Omega_{b}^{-}\to\Omega_{c}^{**0}\rho^{-})$, as listed in Table 3. In particular, we highlight several features of our predictions for the $1P$-wave $\Omega_{c}$ states. First, the predicted branching fractions ${\cal B}(\Omega_{b}^{-}\to\Omega_{c}(3065)^{0}\rho^{-})$ and ${\cal B}(\Omega_{b}^{-}\to\Omega_{c}(3090)^{0}\pi^{-})$ are of order $10^{-2}$, which are about five and eight times larger than their corresponding $\pi$- and $\rho$-mode counterpart channels, respectively. Such significant differences can serve as useful tests of the constituent quark model (CQM) employed in this work. Second, the $1P$-wave $\Omega_{c}$ state $|J^{P}=1/2^{-},0\rangle_{\lambda}$ is denoted by $\Omega^{\prime}_{c}(3000)$ in Table 3, as it is predicted to have a mass close to that of $\Omega_{c}(3000)$. We find that ${\cal B}(\Omega_{b}^{-}\to\Omega^{\prime}_{c}(3000)^{0}\pi^{-})$ is comparable to those of other $1P$-wave states. However, since the decay width of $\Omega^{\prime}_{c}(3000)^{0}$ is expected to be very broad, $\Gamma=80-120$ MeV Zhong:2025oti ; Wang:2017hej , in the resonant process $\Omega_{b}^{-}\to\Omega^{\prime}_{c}(3000)^{0}\pi^{-}$ followed by $\Omega^{\prime}_{c}(3000)^{0}\to\Xi_{c}^{+}K^{-}$, the corresponding resonance may appear broad and flat, making it difficult to observe in the $\Xi_{c}^{+}K^{-}$ invariant mass spectrum.
Apart from the five $\Omega_{c}^{**}$ baryons, the LHCb collaboration reported the two excited $\Omega_{c}^{*}$ states $\Omega_{c}(3185)$ and $\Omega_{c}(3327)$ in the proton-proton collision LHCb:2023sxp , not confirmed in the $e^{+}e^{-}$ collision by Belle yet. This leaves the room for the examinations in $\Omega_{b}$ weak decays. Thus, we obtain the branching fractions of $\Omega_{b}^{-}\to\Omega_{c}(3185)^{0}\pi^{-}$ and $\Omega_{b}^{-}\to\Omega_{c}(3185)^{0}\rho^{-}$ to be $(6.0,1.4)\times 10^{-3}$, respectively. The $\Omega_{c}(3185)$ baryon is commonly interpreted as a $2S$-wave state with quantum numbers $J^{P}=3/2^{+}$ Zhong:2025oti ; Yu:2023bxn ; Karliner:2023okv ; Pan:2023hwt ; Kucukyilmaz:2025rsd , although alternative interpretations have also been proposed Xin:2023gkf ; Yan:2023ttx ; Yan:2023tvl ; Ozdem:2023okg ; Feng:2023ixl .
On the other hand, the theoretical investigations of Refs. Zhong:2025oti ; Wang:2023wii ; Luo:2023sra ; Pan:2023hwt ; Yu:2023bxn ; Garcia-Tecocoatzi:2024aqz ; Li:2024zze ; Weng:2024roa interpret the $\Omega_{c}(3327)$ baryon observed in Ref. LHCb:2023sxp as a candidate $1D$-wave excitation, leaving its spin-parity quantum numbers $J^{P}$ undetermined. Weak decays of the $\Omega_{b}$ baryon, such as $\Omega_{b}^{-}\to\Omega_{c}(3327)^{0}\pi^{-}$ and $\Omega_{b}^{-}\to\Omega_{c}(3327)^{0}\rho^{-}$, may provide a useful probe for identifying the quantum numbers of $\Omega_{c}(3327)^{0}$. In our framework, six $1D$-wave $\Omega_{c}$ states are listed in Table 3. Among them, the states $|J^{P}=5/2^{+},3\rangle_{\lambda}$ and $|J^{P}=7/2^{+},3\rangle_{\lambda}$, whose predicted masses are close to $3327~\text{MeV}$, are plausible candidates for the newly observed resonance. Assuming $\Omega_{c}(3327)$ to be a $1D$-wave excitation with $J^{P}=5/2^{+}$ or $7/2^{+}$ Zhong:2025oti ; Luo:2023sra , we predict
${\cal B}(\Omega_{b}^{-}\to\Omega_{c}(3327)^{0}\pi^{-},\Omega_{c}(3327)^{0}\rho^{-})=(2.0,5.7)\times 10^{-3}~\text{or}~(4.6,0.8)\times 10^{-3}\,.$ (19)
The pronounced differences between these two scenarios provide a clear discriminant that can be tested in future measurements at LHCb. As a final remark, one and four $\Omega_{c}^{*}$ baryons in the $2S$ classification and the $1D$ classification listed in Table 3, respectively, remain unexplored in both scattering processes and weak decays. The branching fractions predicted in this work therefore provide useful guidance for future experimental searches.
In summary, we have carried out a systematic study of the $\Omega_{c}$ spectroscopy up to the $1D$ excitations by analyzing the non-leptonic decays of the $\Omega_{b}^{-}$ baryon within the constituent quark model framework. The predicted branching fractions for $\Omega_{b}^{-}\to\Omega_{c}^{0}\pi^{-}$ and $\Omega_{b}^{-}\to\Omega_{c}^{0}\rho^{-}$, at the level of $10^{-3}$, have been found to be consistent with existing theoretical estimates. Four of the five $\Omega_{c}^{**}$ states, $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$, and $\Omega_{c}(3090)$, have been experimentally observed in $\Omega_{b}$ decays; however, their production rates have not yet been measured. We therefore interpret these states as members of the $1P$-wave multiplet and provide predictions for the branching fractions of $\Omega_{b}^{-}\to\Omega_{c}^{**}\pi^{-},\Omega_{c}^{**}\rho^{-}$ at the order of $10^{-3}$. We have further investigated the newly observed $\Omega_{c}(3327)$ state by considering possible $1D$-wave assignments with $J^{P}=5/2^{+}$ and $7/2^{+}$. The predicted branching fractions for $\Omega_{b}^{-}\to\Omega_{c}(3327)^{0}\pi^{-}$ and $\Omega_{b}^{-}\to\Omega_{c}(3327)^{0}\rho^{-}$ exhibit distinct patterns for the two scenarios, providing a sensitive probe of the spin-parity assignment and internal structure of this state. These predictions can be tested in future measurements at LHCb.
Acknowledgements
KLW was supported by the National Natural Science Foundation of China (Grants No. 12205026), and Support Project for Young Teachers’ Research and Innovation Abilities under Grant No.2025Q037. YKH was supported by the National Natural Science Foundation of China (Grants No. 12575101 and No. 12175128). JW was supported by Shanxi Provincial Graduate Education Innovation Program Project (Grant No. 2025XS111).